OpenStudy (anonymous):

when solving for interest on a annuity. 6000((1+0.06)^25 (-1) / 0.06. could some one explain to me how to do it

OpenStudy (anonymous):

\[6000\frac{ (1+0.06)^{25}-1 }{ 0.06 }\]

OpenStudy (anonymous):

-1 is part of the formula for an annuity

OpenStudy (anonymous):

The rate r = .06/12

OpenStudy (anonymous):

this is an alg equation

OpenStudy (anonymous):

25 years

OpenStudy (anonymous):

rate 0.06

OpenStudy (anonymous):

So t = 25 * 12 = 300 months r = .06/12 = .005 That's how you calculate Future value with annuity formula

OpenStudy (anonymous):

that is not how it was given see above

OpenStudy (anonymous):

ok so what do i do with the two answers

OpenStudy (anonymous):

ok, just very confused with the algb2

OpenStudy (anonymous):

Then the result * ( 6000/ .005) = 4, 157, 963.78

OpenStudy (anonymous):

ok ill try it and see

OpenStudy (anonymous):

can you show me on the equation board, i am still lost

OpenStudy (anonymous):

FV = 6,000 ( 1.005^300 -1 ) / .005